# Diffusion and instabilities in large-N holographic Fermi liquids: the vector fluctuations of the electron star

### v2

@@ 1 Introduction
More specifically, we want to understand the physics of the simplest holographic Fermi liquid model – the electron star – by studying its hydrodynamic response. For all we know, electron stars represent something like multiplets of flavored Lifshitz Fermi liquids. We may indeed call them Fermi liquids as they have a sharp Fermi surface and long-living quasiparticles, but they are certainly not Landau Fermi liquids – the thermodynamic quantities and the self-energy of the quasiparticle scale very differently, and there is a large number (strictly speaking infinity) of Fermi surfaces. So far, works in this direction include the brief consideration of conductivity in the original paper Hartnoll:2010gu and the study of quasinormal modes and hydrodynamics in Kaplis:2013gux ; Gran:2018jnt . In Gran:2018jnt , the authors consider the hydrodynamic response in longitudinal sector but the focus is on modes with Robin boundary conditions encoding for a double-trace deformation corresponding to a plasmon Gran:2017jht ; Gran:2018vdn . Nevertheless, even with standard (Dirichlet) boundary conditions, Gran:2018jnt finds a subdominant contribution from a mode with an anomalous dispersion relation, a hint that some unusual phenomena might be expected. In Kaplis:2013gux , the author studies the diffusive quasinormal mode and finds an anomalous temperature dependence of the diffusion coefficient.
@@
Inspired by these works, we systematically compute the response functions and dispersion relations of the transverse sector (the longitudinal sector was studied to some extent in Kaplis:2013gux ; Gran:2018jnt and will be studied further in a separate work) for various fermion charges and various temperatures, including $T=0$. Our intention is to learn about the true nature of this somewhat pathological large-$N$ system using the hydrodynamic response as a diagnostic tool. While we do not think that the electron star as such is a very realistic model of any condensed matter system, it is important to understand its workings in order to construct a more meaningful model. In hindsight, we will find some big surprises, in particular an instability at intermediate temperatures, suggesting that the simple fluid approximation is a false vacuum in some parameter range.
At this place it is useful to fix some terminology. While the zero-temperature solution is universally called the electron star, the finite-temperature variant was originally called simply the electron star at finite temperature Hartnoll:2010ik but later on, e.g. in GiangrecoMPuletti:2015grq ; Gran:2018jnt it was called the electron cloud, a more illustrative term as it captures the fact that at finite temperature the electron density is only nonzero at some distance from the horizon. We thus exclusively use the term electron cloud for the thermal case and reserve the name electron star solely for the zero-temperature case.
@@ 2.1 The fluid limit and the equations of motion
Here, $\phi$ is the auxiliary (Clebsch) potential and $\lambda$ is a Lagrange multiplier (for details see Hartnoll:2010gu ; we will not make explicit use of this action). In the ideal fluid approximation, the above general fluid action further simplifies to just the pressure of the fluid when calculated on-shell. The summation over the Fermi sea (i.e., over the occupied levels of the fermions) now becomes the integration, yielding closed-form expressions for densities and pressure: [^1] In this approximation the results for $\rho$, $\sigma$ and $p$ are not sensitive to curvature and have the same form as in flat space Hartnoll:2010gu .
$\hat{\rho}=\hat{\beta}\int_{\hat{m}}^{\frac{h}{\sqrt{f}}}\epsilon^{2}\sqrt{\epsilon^{2}-\hat{m}^{2}}\mathrm{d}\epsilon,\quad\hat{\sigma}=\hat{\beta}\int_{\hat{m}}^{\frac{h}{\sqrt{f}}}\epsilon\sqrt{\epsilon^{2}-\hat{m}^{2}}\mathrm{d}\epsilon,\quad-\hat{p}=\hat{\rho}-\frac{h}{\sqrt{f}}\hat{\sigma},$ (6)
@@
$\hat{p}=L^{2}\kappa^{2}p,\quad\hat{\rho}=L^{2}\kappa^{2}\rho,\quad\hat{\sigma}=eL^{2}\kappa\sigma,\quad\hat{\beta}=\frac{e^{4}L^{2}}{\pi^{2}\kappa^{2}},\quad\hat{m}^{2}=\frac{\kappa^{2}}{e^{2}}m^{2}.$ (7)
We thus have two microscopic free parameters to vary: the fermion mass $\hat{m}$ and the level density constant of the fermion $\hat{\beta}$. The fermion mass must satisfy the inequality $10\leq\hat{m}<$ 1 Hartnoll:2010gu , while $\hat{\beta}$ is in principle arbitrary, but cannot be too large if the fluid approximation is to be valid Leiden:2011 . Taking into account the requirement of string theory that the gravitational coupling equals the square of the Maxwell coupling, we may assume that in the classical gravity regime:
@@
which means that $\hat{\beta}\sim 1$ Hartnoll:2010gu . For further use, we define also the local bulk chemical potential as Hartnoll:2010gu
$\mu_{\mathrm{loc}}=\frac{A_{t}}{L\sqrt{f}}=\frac{e}{\kappa}\frac{h}{\sqrt{f}}.$ (9)
@@
This ansatz encapsulates the most general radially symmetric static solution, homogeneous and isotropic along the transverse coordinates $(x,y)$, with no magnetic field or stationary current. The AdS boundary is located at $r=0$. Now, finally, from Eqs. (1) and (5) and the above ansatz, the equations of motion are found to be Hartnoll:2010gu :
@@
$h^{\prime\prime}+\frac{g\hat{\sigma}}{\sqrt{f}}\left(\frac{rhh^{\prime}}{2}-f\right)$ $=$ $0$ (14)
@@ 2.2.2 Finite-temperature solutions: electron clouds
The electron cloud exists for $0<T<T_{c}$, where $T_{c}$ is some critical temperature. At $T_{c}$ the cloud vanishes and the system becomes a fully fractionalized semilocal quantum liquid Vegh:2009 ; Faulkner:2009wj ; Liu:2011 . At the critical temperature we have $r_{1}=r_{2}\equiv r_{c}$. It turns out Hartnoll:2010ik that the critical radius and the critical temperature are given by the condition that the cloud contains just a single fermion, located at the maximum of the bulk chemical potential. This yields the equations for $r_{c}$ and $T_{c}$:
@@ 3.1 Fluctuation equations
$a_{y}^{\prime\prime}+\frac{1}{2}\left(\frac{f^{\prime}}{f}-\frac{g^{\prime}}{g}\right)a_{y}^{\prime}-g\left(k^{2}r^{2}-\frac{\omega^{2}}{f}\right)a_{y}+\frac{h^{\prime}}{r^{2}f}h^{y\prime}_{\phantom{y}t}+\frac{g\hat{\sigma}}{r^{2}\sqrt{f}-r}h^{y}_{\phantom{y}t}$ $=$ $0$ (26)
@@
$02r^{2}h^{\prime}a_{y}+h^{y\prime}_{\phantom{y}t}+\frac{kr^{2}f}{\omega}h^{x\prime}_{\phantom{x}y}=$ 0 . (27)
@@
$\left(\sqrt{\frac{f}{g}}Y^{\prime}\right)^{\prime}-\left(\left(k^{2}r^{2}f-\omega^{2}\right)\sqrt{\frac{g}{f}}+\frac{2{h^{\prime}}^{2}}{\sqrt{fg}}+\frac{f\sqrt{g}\hat{\sigma}}{h}\right)Y+kr^{2}h^{\prime}Z$ $=$ $0$ (34)
@@ 4 Results: correlation functions and hydrodynamics
Apart from the single exception of $z=20$ at $T/T_{c}=0.05$, [^7] This case exhibits very peculiar features and generally does not follow any of the trends common to the rest of the solutions. It might be that such behavior emerges when one approaches the zero temperature in general, but for now we are not able to give any definite answer to this puzzle. similar results are obtained for all three $z$ values so we choose to present the results for $z=4$. Since all the correlators depending on $G_{+}$ exhibit similar properties, we present only $G_{ytyt}$. Figures 3-5 show its frequency dependence for various momenta and five different temperatures: $T/T_{c}=0$, $T/T_{c}=0.23$ (Figure 3), $T/T_{c}=0.35$, $T/T_{c}=0.65$ (Figure 4) and $T/T_{c}=0.95$ (Figure 5). [^8] The imaginary parts are given with a negative sign since $-\mathrm{Im}\,G_{ytyt}$ is always positive. The bottom row in Figure 5 is again for $T/T_{c}=0.95$ but for the RN black hole for comparison. The corresponding $k$-$\omega$ density plots are shown in Figures 6-8 (for these figures we sometimes use denser momentum grids in order to show the full 2D energy-momentum dependence). We express $\omega$ and $k$ in the units of the field theory chemical potential. [^9] There are two scales in our theory: the field theory chemical potential $\hat{\mu}$ and the temperature $T$. We have found the former more convenient to work with.
@@
    At intermediate temperatures (Figures 4 and 7) the correlation functions look very differently and clearly cannot be described by (normal) hydrodynamics.
[…]
@@
At intermediate temperatures, both subsystems (the horizon and the cloud) are non-negligible and their interplay drastically changes the picture. It looks as if the diffusion pole moves towards higher frequencies and momenta as the temperature increases (until it increases so much that we re-enter the hydrodynamic regime). For now, this is just a handwaving claim based on the visual inspection of the figures; in Section 5.1 we will corroborate this analytically. In any case, there is no sign of a hydrodynamic regime at intermediate $T/T_{c}$. In fact, things are even more bizarre at intermediate temperatures. In the next section we will see that this state is unstable – a pole develops in the upper half of the complex frequency plane, signaling that the electron cloud solution might be the false vacuum in this regime.
As a further check on the existence of the hydrodynamic regime, let us assume a dispersion relation of the form $\omega\propto k^{\gamma}$ ($\gamma\in\mathbb{R}$) and try to determine $\gamma$ by fitting. For given $z$ and $T$, we select a couple of lowest momenta $k_{i}$ (roughly a third of the total momentum grid) and determine for each of them the position $\omega_{i}$ of the local maximum of $\mathrm{Im}\,G_{+}$. The dispersion relation is then estimated from the linear fit of the log-log plot of the data $(k_{i},\omega_{i})$. We show two examples in Figure 9 – one giving the expected result for diffusion, i.e. $\gamma\sim 2$, and the other giving a different power law ($\gamma\sim 1$). The results are summarized in Table 1. Since we only fit a small number of points, this is just a rough estimate. Nevertheless, it is illustrative enough and in agreement with our observations. What is more, it suggests a possible linear dispersion regime (i.e. $\omega\propto k$) at intermediate temperatures. This will prove to be true, and we will see in Section 5.1 how it is responsible for the emergence of poles in the upper half-plane.
Figure 9: The maxima of $\mathrm{Im}\,G_{+}$ for $z=4$, for $T/T_{c}=0$ (left) and $T/T_{c}=0.65$ (right). In the first case the fitting gives $\gamma\approx 2.2$, in rough agreement with the expected quadratic scaling. In the second case $\gamma\approx 1.3$, in rough agreement with our conclusion that at the intermediate temperatures the hydrodynamics vanishes. The line is just to guide the eye.
$T/T_{c}$ 0 0.05 0.23 0.35 0.65 0.95
$z=2.5$ 2.3 - - 1.9 1.1 2.5
$z=4$ 2.2 - 0.9 1.1 1.3 2.8
$z=20$ 1.7 1.8 2.0 2.4
Table 1: The approximate values of the exponent $\gamma$ obtained by fitting.
@@
Transverse conductivity $\hat{\sigma}_{y}$ is given in Figure 10. It has a similar profile at all temperatures and for all Lifshitz exponents. The only difference is the value of $\mathrm{Re}\,\hat{\sigma}_{y}$ in the limit $\omega\to 0$, which grows with temperature for every $z$ [^11] It is understood that extending $\mathrm{Re}\,\hat{\sigma}_{y}$ to $\omega=0$ gives a $\delta$-function. Otherwise, the conductivity is mainly featureless. Specifically, there is no sign of a diffusion or any other pole at low frequencies, so we infer that $G_{-}$ (see Eq. (51)) is a holomorphic function in the entire complex frequency plane. In Section 5.1 we will confirm analytically that this is true at least in the low-energy limit.
@@ 5 Semianalytical low-energy expansion
and where all the derivatives are with respect to $r_{*}$, while $r$ is understood to be a funciton of $r_{*}$.
@@
[…]
@@
Equations (65-67) can also be easily solved provided that we know the two independent solutions – let us call them $\upsilon_{0}^{(1)}$ and $\upsilon_{0}^{(2)}$ – to Eq. (65), since the homogeneous parts always have the same form. However, as we mentioned above, these solutions can only be found numerically. In order to distinguish the two numerical solutions, we set independent initial conditions at the horizon. We find:
@@
where $W_{1,2}$ is the Wronskian of $\upsilon_{0}^{(1)}$ and $\upsilon_{0}^{(2)}$.

### v3

@@ 1 Introduction
More specifically, we want to understand the physics of the simplest holographic Fermi liquid model – the electron star – by studying its hydrodynamic response. For all we know, electron stars represent something like multiplets of flavored Lifshitz Fermi liquids. We may indeed call them Fermi liquids as they have a sharp Fermi surface and long-living quasiparticles, but they are certainly not Landau Fermi liquids – the thermodynamic quantities and the self-energy of the quasiparticle scale very differently, and there is a large number (strictly speaking infinity) of Fermi surfaces. So far, works in this direction include the brief consideration of conductivity in the original paper Hartnoll:2010gu and the study of quasinormal modes and hydrodynamics in Kaplis:2013gux ; Gran:2018jnt . In Gran:2018jnt , the authors consider the hydrodynamic response in the longitudinal sector but the focus is on modes with Robin boundary conditions encoding for a double-trace deformation corresponding to a plasmon Gran:2017jht ; Gran:2018vdn . Nevertheless, even with standard (Dirichlet) boundary conditions, Gran:2018jnt finds some modes with exotic dispersion relations (non-diffusive scaling, modes hitting the real axis at specific momenta, etc), a hint that some unusual phenomena might be expected. In Kaplis:2013gux , the author studies the diffusive quasinormal mode and finds an anomalous temperature dependence of the diffusion coefficient.
@@
Inspired by these works, we systematically compute the response functions and dispersion relations of the transverse sector (as we mention above, the longitudinal sector was studied in Kaplis:2013gux ; Gran:2018jnt and will be studied further in a separate work) for various fermion charges and various temperatures, including $T=0$. Our intention is to learn about the true nature of this somewhat pathological large-$N$ system using the hydrodynamic response as a diagnostic tool. While we do not think that the electron star as such is a very realistic model of any condensed matter system, it is important to understand its workings in order to construct a more meaningful model. In hindsight, we will find some big surprises, in particular an instability at intermediate temperatures, suggesting that the simple fluid approximation is a false vacuum in some parameter range.
At this place it is useful to fix some terminology. The zero-temperature solution is universally called the electron star. The finite-temperature variant was originally called simply the electron star at finite temperature Hartnoll:2010ik but later on, e.g. in GiangrecoMPuletti:2015grq ; Gran:2018jnt , it was called the electron cloud, a more illustrative term as it captures the fact that at finite temperature the electron density is only nonzero at some distance from the horizon. We thus exclusively use the term electron cloud for the thermal case and reserve the name electron star solely for the zero-temperature case.
@@ 2.1 The fluid limit and the equations of motion
Here, $\phi$ is the auxiliary (Clebsch) potential and $\lambda$ is a Lagrange multiplier (for details see Hartnoll:2010gu ; we will not make explicit use of this action). In the ideal fluid approximation, the above general fluid action further simplifies to just the pressure $p$ of the fluid when calculated on-shell. The summation over the Fermi sea (i.e., over the occupied levels of the fermions) now becomes the integration, yielding closed-form expressions for densities and pressure: [^1] In this approximation the results for $\rho$, $\sigma$ and $p$ are not sensitive to curvature and have the same form as in flat space Hartnoll:2010gu .
$\hat{\rho}=\hat{\beta}\int_{\hat{m}}^{\hat{\mu}_{\mathrm{loc}}}\epsilon^{2}\sqrt{\epsilon^{2}-\hat{m}^{2}}\mathrm{d}\epsilon,\quad\hat{\sigma}=\hat{\beta}\int_{\hat{m}}^{\hat{\mu}_{\mathrm{loc}}}\epsilon\sqrt{\epsilon^{2}-\hat{m}^{2}}\mathrm{d}\epsilon,\quad-\hat{p}=\hat{\rho}-\hat{\mu}_{\mathrm{loc}}\hat{\sigma}$, (6)
@@
$\hat{p}=L^{2}\kappa^{2}p,\quad\hat{\rho}=L^{2}\kappa^{2}\rho,\quad\hat{\sigma}=eL^{2}\kappa\sigma,\quad\hat{\beta}=\frac{e^{4}L^{2}}{\kappa^{2}}\beta,\quad\hat{m}^{2}=\frac{\kappa^{2}}{e^{2}}m^{2},\quad\hat{\mu}_{\mathrm{loc}}=\frac{\kappa}{e}\mu_{\mathrm{loc}},$ (7)
and where we define the local bulk chemical potential as Hartnoll:2010gu
$\mu_{\mathrm{loc}}=\frac{A_{t}}{\sqrt{-g_{tt}}}.$ (8)
We thus have two microscopic free parameters to vary: the fermion mass $\hat{m}$ and the level density constant of the fermion $\hat{\beta}$. The fermion mass must satisfy the inequality $0\leq\hat{m}<1$ Hartnoll:2010gu , while $\hat{\beta}$ is in principle arbitrary, but cannot be too large if the fluid approximation is to be valid Leiden:2011 . Taking into account the requirement of string theory that the gravitational coupling equals the square of the Maxwell coupling, we may assume that in the classical gravity regime:
@@
which means that $\hat{\beta}\sim 1$ Hartnoll:2010gu .
@@
This ansatz encapsulates the most general radially symmetric static solution, homogeneous and isotropic along the transverse coordinates $(x,y)$, with no magnetic field or stationary current. The AdS boundary is located at $r=0$. Now, finally, from Eqs. (1), (5) and (6-8) and the above ansatz, the equations of motion are found to be Hartnoll:2010gu :
@@
$h^{\prime\prime}+\frac{g\hat{\sigma}}{\sqrt{f}}\left(\frac{rhh^{\prime}}{2}-f\right)$ $=$ $0.$ (14)
@@ 2.2.2 Finite-temperature solutions: electron clouds
The electron cloud exists for $0<T<T_{c}$, where $T_{c}$ is some critical temperature. At $T_{c}$ the cloud vanishes and only the RN black hole remains. This means that for $T>T_{c}$ all the charge is distributed over the horizon. The holographic dual of this phase is interpreted as the fully fractionalized semilocal quantum liquid, conjectured to describe a strange metal Vegh:2009 ; Faulkner:2009wj ; Liu:2011 .
At the critical temperature we have $r_{1}=r_{2}\equiv r_{c}$. It turns out Hartnoll:2010ik that the critical radius and the critical temperature are given by the condition that the cloud contains just a single fermion, located at the maximum of the bulk chemical potential. This yields the equations for $r_{c}$ and $T_{c}$:
@@ 3.1 Fluctuation equations
$a_{y}^{\prime\prime}+\frac{1}{2}\left(\frac{f^{\prime}}{f}-\frac{g^{\prime}}{g}\right)a_{y}^{\prime}-g\left(k^{2}r^{2}-\frac{\omega^{2}}{f}\right)a_{y}+\frac{h^{\prime}}{r^{2}f}h^{y\prime}_{\phantom{y}t}+\frac{g\hat{\sigma}}{r^{2}\sqrt{f}-r}h^{y}_{\phantom{y}t}$ $=$ $0.$ (26)
@@
$2r^{2}h^{\prime}a_{y}+h^{y\prime}_{\phantom{y}t}+\frac{kr^{2}f}{\omega}h^{x\prime}_{\phantom{x}y}=0.$ (27)
@@
$\left(\sqrt{\frac{f}{g}}Y^{\prime}\right)^{\prime}-\left(\left(k^{2}r^{2}f-\omega^{2}\right)\sqrt{\frac{g}{f}}+\frac{2{h^{\prime}}^{2}}{\sqrt{fg}}+\frac{f\sqrt{g}\hat{\sigma}}{h}\right)Y+kr^{2}h^{\prime}Z$ $=$ $0.$ (34)
@@ 4 Results: correlation functions and hydrodynamics
Apart from the single exception of $z=20$ at $T/T_{c}=0.05$, [^7] This case exhibits very peculiar features and generally does not follow any of the trends common to the rest of the solutions. It might be that such behavior emerges when one approaches the zero temperature in general, but for now we are not able to give any definite answer to this puzzle. similar results are obtained for all three $z$ values so we choose to present the results for $z=4$. Since all the correlators depending on $G_{+}$ exhibit similar properties, we present only $G_{ytyt}$. Figures 3-5 show its frequency dependence for various momenta and five different temperatures: $T_{c}=0, T/T$$T_{c}=0.23$ (Figure 3), $T_{c}=0.35, T/T$$T_{c}=0.65$ (Figure 4) and $T/T_{c}=0.95$ (Figure 5). [^8] The imaginary parts are given with a negative sign since $\mathrm{Im}\,G_{ytyt}$ is always positive. The bottom row in Figure 5 is again for $T/T_{c}=0.95$ but for the RN black hole for comparison. The corresponding $k$-$\omega$ density plots are shown in Figures 6-8 (for these figures we sometimes use denser momentum grids in order to show the full 2D energy-momentum dependence). We express $\omega$ and $k$ in the units of the field theory chemical potential. [^9] There are two scales in our theory: the field theory chemical potential $\hat{\mu}$ and the temperature $T$. We have found the former more convenient to work with.
@@
    At intermediate temperatures (Figures 4 and 7) the correlation functions look very different and clearly cannot be described by (normal) hydrodynamics.
It is not that difficult to understand the above points on a qualitative level. At low temperatures, the horizon is small (at $T=0$ it is nonexistent) and the response of the system is dominated by the electron star. The fact that it is a fluid system, that it has metallic conductivity as found in Hartnoll:2010gu and that it is approximately described by Lifshitz geometry in IR, which is known to exhibit hydrodynamic behavior at finite (but arbitrarily small) temperatures at least for $z>2$ Sybesma:2015oha ; Gursoy:2016tgf all suggest that it could have a hydrodynamic response. At high temperatures the horizon is large and the electron cloud is small, hence the system should not deviate much from the RN hydrodynamics. The RN hydrodynamics is very well studied Davison:2013bxa , and this is why we make a comparison between the electron cloud and the pure RN black hole at $T/T_{c}=0.95$. The comparison is handily eased by the fact that in the pure RN case $\hat{\mu}=\hat{\mu}_{0}\approx 0.65$, while in the presence of the cloud $\hat{\mu}\approx 0.67$. [^10] Another way to compare the two geometries is to recast Eq. (34) into a Schrödinger form by setting $Z=0$ and analyze the resulting effective potential. The effective potentials in the two cases agree in the asymptotic regions (in the IR they are the same), but may differ considerably in the region occupied by the electron cloud. In general, the effective potential in the presence of the cloud is deeper and admits more bound states, which agrees with the core definition of the electron star.
@@
At intermediate temperatures, both subsystems (the horizon and the cloud) are non-negligible and their interplay drastically changes the picture. It looks as if the diffusion pole moves towards higher frequencies and momenta as the temperature increases (until it increases so much that we re-enter the hydrodynamic regime). For now, this is just a handwaving claim based on the visual inspection of the figures; in Section 5.1 we will corroborate this analytically. In any case, there is no sign of a hydrodynamic regime at intermediate $T/T_{c}$. In fact, things are even more bizarre at intermediate temperatures. In the next section we will see that this state is unstable – a pole develops in the upper half of the complex frequency plane, signaling that the electron cloud solution might be a false vacuum in this regime.
@@
Transverse conductivity $\hat{\sigma}_{y}$ is given in Figure 9. It has a similar profile at all temperatures and for all Lifshitz exponents. The only difference is the value of $\mathrm{Re}\,\hat{\sigma}_{y}$ in the limit $\omega\to 0$, which grows with temperature for every zz. [^11] It is understood that extending $\mathrm{Re}\,\hat{\sigma}_{y}$ to $\omega=0$ gives a $\delta$-function. Otherwise, the conductivity is mainly featureless. Specifically, there is no sign of a diffusion or any other pole at low frequencies, so we infer that $G_{-}$ (see Eq. (51)) is a holomorphic function in the entire complex frequency plane. In Section 5.1 we will confirm analytically that this is true at least in the low-energy limit.
@@ 5 Semianalytical low-energy expansion
and where all the derivatives are with respect to $r_{*}$, while $r$ is understood to be a function of $r_{*}$.
@@
with $\zeta_{n}\sim\upsilon_{n}\sim\epsilon^{n}$ for an integer $n\geq 0$. We assume that $k\sim\epsilon$ and $\omega\sim\epsilon^{\gamma}$ for some real parameter $\gamma$, $\gamma\geq 1$. In this way we allow that $\omega\sim|k|^{\gamma}$ need not satisfy the normal diffusive dispersion relation. However, this immediately poses a difficulty: because of the scaling $\omega\sim\epsilon^{\gamma}$ we cannot know in advance at which order in the $\epsilon$-expansion the $\omega$-dependent terms should appear. These terms may be of order $\epsilon$, $\epsilon^{2}$, or somewhere “in-between” – $\gamma$ is not known in advance and need not be an integer. But this will turn out irrelevant as we will eventually see, since the structure of the equations allows the leading $\omega$-term to be captured. Thus, we expand the equations assuming first that $\omega\sim\epsilon$. Then to second order we get:
@@
Equations (65-67) can also be easily solved provided that we know two independent solutions to Eq. (65) (let us call them $\upsilon_{0}^{(1)}$ and $\upsilon_{0}^{(2)}$), since the homogeneous parts of all three equations have the same form. However, as we mentioned above, these solutions can only be found numerically. In order to distinguish the two numerical solutions, we set independent initial conditions at the horizon. We find:
@@
where $W_{1,2}$ is the Wronskian of $\upsilon_{0}^{(1)}$ and $\upsilon_{0}^{(2)}$. Now that we have the full solution, we can come back to our comment after Eqs. (60-61) about the difficulties arising from the fact that $\gamma$ is not known in advance. The difficulties are resolved by the fact that all inhomogeneous equations contain a pure $\omega$-term (i.e., without multiplication by $k$), so whatever the scaling of $\omega$ with $k$, the leading terms with $\omega$ will be included at every order in $k$.